The physical fitness of an athlete is often measured by how much oxygen the athlete takes in (which is recorded in milliliters per kilogram, ml/kg). The mean maximum oxygen uptake for elite athletes has been found to be 70 with a standard deviation of 8.2. Assume that the distribution is approximately normal.

(a)  What is the probability that an elite athlete has a maximum oxygen uptake of at least 55 ml/kg?

answer:

(b)  What is the probability that an elite athlete has a maximum oxygen uptake of 60 ml/kg or lower?

answer:

(c)  Consider someone with a maximum oxygen uptake of 36 ml/kg. Is it likely that this person is an elite athlete? Write "YES" or "NO."
Answer #1

P(X < A) = P(Z < (A - mean)/standard deviation)

Mean = 70 ml/kg

Standard deviation = 8.2 ml/kg

a) P(at least 55 ml/kg) = P(X 55)

= 1 - P(X < 55)

= 1 - P(Z < (55 - 70)/8.2)

= 1 - P(Z < -1.83)

= 1 - 0.0336

= 0.9664

b) P(60 ml/kg or lower) = P(X 60)

= P(Z < (60 - 70)/8.2)

= P(Z < -1.22)

= 0.1112

c) P(36 ml/kg or less) = P(X 36)

= P(Z < (36 - 70)/8.2)

= P(Z < -4.15)

= 0

The probability that maximum oxygen uptake is 36 ml/kg is almost 0. So, this is a highly unlikely event (any event with probability less than 0.05 can be considered as an unlikely event). Therefore, this person must not be an elite athlete.

Ans: NO